OpenStudy (mony01):

Determine whether the equation is separable or not. Write separable ones in separable form. Determine constant solutions, if any. y'=1+y

OpenStudy (loser66):

It is

OpenStudy (loser66):

To write it separable, you just write \(y' = \dfrac{dy}{dx}\) then let y in the left hand side and x in the right hand side

OpenStudy (loser66):

knock knock!! do you get me??

zepdrix (zepdrix):

who's there? :o

OpenStudy (mony01):

not really

OpenStudy (loser66):

@zepdrix , Honorary Professor of Mathematics, help pleaaaaaasssse.

zepdrix (zepdrix):

\[\Large\rm \frac{dy}{dx}=(1+y)\]"Multiply" dx to the other side,\[\Large\rm dy=(1+y)dx\]So from here, how do we get all the y stuff to one side? Hmmm.

zepdrix (zepdrix):

What do you think cash mony?

OpenStudy (mony01):

integrate?

zepdrix (zepdrix):

No, you can't integrate (1+y) against dx.

OpenStudy (mony01):

subtract y?

OpenStudy (mony01):

divide (1+y)

zepdrix (zepdrix):

yes, divide the (1+y) to the other side.

zepdrix (zepdrix):

\[\Large\rm \frac{dy}{1+y}=dx\]Now you have your x's and y's separated. You can integrate at this point.

OpenStudy (mony01):

is it \[\ln \left| 1+y \right|\]

OpenStudy (mony01):

how to you integrate? @zepdrix

zepdrix (zepdrix):

\[\Large\rm \int\limits \frac{dy}{1+y}=\int\limits dx\]Yes. \[\Large\rm \ln|1+y|=x+c\]

OpenStudy (mony01):

is that the answer?

zepdrix (zepdrix):

Determine `constant` solutions? Hmm.

zepdrix (zepdrix):

`Constant` solutions you can equate to `equilibrium` solutions. So this occurs when the first derivative is zero.\[\Large\rm y'=1+y\]\[\Large\rm 0~=1+y\]Solve for y. Does this y value work in our solution that we found?\[\Large\rm \ln|1+y|=x+c\]

OpenStudy (mony01):

y would be -1